# FALCONEye: Finding Answers and Localizing Content in ONE-hour-long videos with multi-modal LLMs

### v2

@@ 4 FALCON-Bench
The videos of our benchmark were sourced from SoccerNet [giancola2018soccernet], MovieChat-1K [song2024moviechat] and Walking Tours Dataset [venkataramanan2023imagenet], ensuring diversity and comprehensive coverage across different video types. Overall, the benchmark comprises 575 questions, covering 4 categories, over 90 videos, with an average video duration of 78.9 minutes and answers localized within a GT temporal window of 38.4 seconds. The dataset is split into a test set (506 questions) and a validation set (70 questions). Each question is accompanied by four different answer options, allowing for multiple-choice question (MCQs) evaluation. However, the primary focus remains on open-ended questions (OQs), as this approach is more aligned with the VAS task. The annotation process was conducted by humans assisted by GPT-4o for generating answer options. More details on the benchmark content and construction process are in Sec. 8 of the supplementary material. The benchmark has been built under the lmms-eval framework [zhang2024lmmseval] to ease further evaluation and will be publicly available upon acceptance.
@@
mPLUG-Owl3 [ye2024mplug] 7B 16 5.52 41.1 28.5 36.0 35.2 0.00 6.63 8.82 4.23 10.0 7.68 0.68 0.29 0.66 0.54 0.00
LLaVA-v1.5 [liu2023llava] 7B 32 11.4 23.5 31.9 44.0 33.1 0.00 12.7 3.92 4.80 2.00 3.57 0.36 0.31 0.24 0.30 0.00
LLaVA-v1.6 [llava-next] 7B 32 12.9 28.4 30.5 38.0 32.3 0.00 14.3 4.90 4.51 2.00 3.80 0.31 0.30 0.20 0.27 0.00
LLaVA-OV [llava-onevision] 7B 32 16.9 48.0 30.2 46.0 41.4 0.00 7.29 9.80 4.23 16.0 10.0 0.53 0.25 1.00 0.59 0.00
Qwen2.5-VL [Qwen2.5-VL] 7B 768 68.9 50.0 24.4 48.0 40.8 0.00 70.1 11.7 8.19 10.0 9.96 0.72 0.53 0.62 0.62 0.00
@@
Video-LLaVA [lin2023video] 7B 8 3.37 28.4 26.8 34.0 29.7 0.00 4.50 6.86 5.08 8.00 6.64 0.56 0.37 0.68 0.53 0.00
VideoChat2-HD [li2024mvbench] 7B 8 6.74 29.4 24.2 26.0 26.5 0.00 8.06 8.82 2.82 8.00 6.54 0.71 0.20 0.52 0.47 0.00
LLaVA-Video [zhang2024llavanext-video] 7B 32 12.4 50.0 38.9 58.0 48.9 0.00 13.0 10.7 11.0 12.0 11.2 0.76 0.66 0.86 0.76 0.00
@@
MovieChat-OV [song2024moviechat] 7B 1024 62.3 43.1 38.7 46.0 42.6 0.00 69.3 9.80 4.80 10.0 8.20 0.53 0.25 0.66 0.48 0.00
Apollo [zohar2024apollo] 7B 0.5fps 36.2 51.9 35.8 64.0 50.5 0.00 36.5 16.6 10.7 18.0 15.1 0.96 0.68 1.06 0.90 0.00
Video-XL-Pro [liu2025video] 3B 256 48.8 52.9 33.8 50.0 45.6 0.00 7B 17.6 7.62 14.0 13.0 1.08 0.45 0.94 0.63 0.00
LLaVA-OV+ReKV [di2025streaming] 7B 0.5fps 92.1 65.6 40.1 56.0 53.9 0.00 96.6 34.3 13.8 24.0 24.0 1.78 0.80 1.40 1.32 0.00
@@
Socratic [zeng2022socratic] 7B+N/A 0.5fps 186.6 53.9 48.0 54.0 51.9 21.3 186.7 25.4 29.0 20.0 24.8 1.49 1.71 1.16 1.45 19.6
LifelongMemory[wang2024lifelongmemory] 7B+N/A 0.5fps 207.3 45.0 41.5 38.0 41.8 0.00 208.2 6.86 22.3 8.00 12.3 0.53 1.37 0.56 0.82 0.00
VideoAgent[wang2024videoagent]+OQ 7B+N/A 9.2 45.0 48.0 52.2 36.0 45.4 19.1 43.8 8.82 24.8 8.00 13.8 0.56 1.38 0.44 0.79 13.3
@@ 5 Experiments
In MCQs, we report accuracy as the primary evaluation metric. For OQs, following [maaz2023video], also applied in others, as InfiniBench [ataallah2024infinibench], we employ GPT-assisted evaluation. GPT reports accuracy (100 or 0) and a score value (0-5 range), assessing key aspects (Correctness of Information, Detailed Orientation, Contextual Understanding, Temporal Understanding and Consistency). We also use Ground Truth over Union (GToU) metric to evaluate the quality of the predicted temporal interval, as visual evidence (short clip) that contains the answer. Section 8.3 of the Supplementary discusses why we find GToU more adequate than mIoU (a clip that tries to match the entire GT clip).
@@
Socratic [zeng2022socratic] 1 13.5K 2.1 12.6 768 14.6 186.7 24.8 1.45 19.6
Socratic-Short [zeng2022socratic] 1 2.4K 1.1 12.6 64 9.8 125.0 15.5 1.03 18.3
LifelongMemory[wang2024lifelongmemory] 10.2 0.3K 1.0 12.6 64 9.8 208.2 12.3 0.82 0.00
VideoAgent[wang2024videoagent]+OQ 5.6 1.3K 3.2 3.70 64 2.2 43.8 13.8 0.79 13.3
@@
A widely used baseline in recent long VQA benchmarks [chandrasegaran2024hourvideo] is the Socratic method [zeng2022socratic]. In this baseline, a VLM generates captions $\mathcal{C}=\{c_{1},\ldots,c_{n}\}$ from short video clips $\{v_{1},\ldots,v_{n}\}\in\mathcal{V}$ and a LLM generates $\mathcal{A}$ from $Q$ and $\mathcal{C}$. In addition to this baseline, we introduce three alternative exploration baselines (Fig. 7 of the supp. material) that rely solely on a VLM, guided by its answer confidence computed as described in Sec 3.
@@
As discussed in the related work, two relevant and recent video agents are VideoAgent [wang2024videoagent] and VideoTree [wang2025videotree]. We use VideoAgent as our reference meta‑architecture to compare with since it couples CLIP and a 7B‑parameter VLM, both of which fit comfortably on a single RTX3090. In contrast, VideoTree depends on EVA‑CLIP 8B—too large to co‑load with a VLM‑7B on that GPU—and its public codebase lacks the VLM captioning module, preventing a fair, fully reproducible baseline.
@@
We evaluate our meta-architecture using Qwen2.5-VL [Qwen2.5-VL] as the VLM and GPT4o-mini [openai2023chatgpt] as the LLM. As detailed in Sec. 6, Qwen2.5-VL outperforms other small-size VLMs on FALCON-Bench OQs in short clips near the answer, and is well calibrated. Its ViT supports images of any resolution, dynamically adjusting visual tokens, enabling an accurate zoom-in effect as clip lengths decrease. GPT4o-mini provides a balanced trade-off between cost and reasoning power (Sec. 6). Prompts used for LLM queries are detailed in Sec. 9.2. Inspired by the Socratic approach [chandrasegaran2024hourvideo], first-level clips are 1min length, but captions are now shorter formed by up to 64 tokens (instead of 768). Within these clips, second-order (20s) and third-order (5s) clips are explored. The zoom-in effect, validated in supplementary Sec. 11, adjusts from 30 frames ($672\times 1204$) in first-level clips to 10 frames ($824\times 1462$) in third-order clips. In OQs, the LLM determines the final answer based on completion, confidence ($0.8$ threshold), captions, and temporal localization (as explained in Sec. 3). In MCQs, this evaluation is replaced by a simple confidence-based check with $0.9$ threshold since all possible answers are predefined (A, B, C, or D) and inherently valid, making confidence the only selection criterion. Lastly, we evaluate two versions of FALCONEye: FALCONEye-Pro, the top-performing configuration, which evaluates up to 45 candidate clips, and FALCONEye-Flash, a cost-efficient variant that limits evaluation to 10 candidates.
@@ 5.1 FALCON-Bench results
Table 4 presents the performance of baselines, open-source VLMs, and meta-architectures on FALCON-Bench. As baselines, we include human performance both to establish an upper bound and to analyze human strategies in VAS, which guided our algorithm design (Sec. 8.4). Additionally, we evaluate our two selected LLMs (GPT4o-mini and Qwen2.5-VL) in blind mode (without visual information) to assess their reliance on text-based priors. For open-source VLMs, we test small-scale versions of the main state-of-the-art models, including Apollo and MovieChat, which are specifically designed for long videos. We also evaluate three video agents, Socratic [zeng2022socratic], LifelongMemory [wang2024lifelongmemory] for VQA and VideoAgent [wang2024videoagent]. To ensure a fair comparison with our FALCONEye, these meta-architectures are implemented using the same VLM (Qwen2.5-VL) and LLM (GPT4o-mini). The Socratic approach generates captions (768 tokens) from 1-minute clips, while VideoAgent is extended to handle open-ended questions and predict answer localization, resulting in VideoAgent [wang2024videoagent]++OQ. Most VLMs that are limited to a small number of sampled frames fail to outperform the LLM-blind baselines in MCQs, suggesting that their advantage over random guessing stems from discarding certain options, rather than actual answer retrieval from visual content. A similar issue arises with Qwen2.5-VL when operating in its default mode (2fps, up to 768 frames), as it downscales frames to $140\times 240$, losing crucial details. Only LLaVA-Video, Apollo, and ReKV show clear improvements over the blind baselines. This trend is not observed in OQs, where models must generate answers independently, validating their reliability. Among meta-architectures, only FALCONEye demonstrates robustness for long-form VAS, achieving 70.0% accuracy in MCQs and 44.7% in OQs with its top-performance configuration, and 64.7% and 41.1%, respectively, in the cost-efficient variant. The latter configuration achieves performance close to the Socratic baseline while maintaining a processing time per question not so far from human response times. We analyze their costs and performances, including exploration baselines, in Table 3. We also include Socratic-Short, which utilizes the short captions (64 tokens) generated by FALCONEye. While baseline exploration algorithms outperform existing meta-architectures, they require significantly more time. Notably, our LLM-guided exploration reduces this overhead while improving performance by efficiently narrowing the search.
@@ 5.2 FALCONEye generalization evaluation
To explore FALCONEye generalization beyond our benchmark, we evaluate it out-of-the-box on a standard VQA setting: shorter videos and a broader mix of question types beyond single-detail (SD). With this goal, we select MLVU [zhou2024mlvu], the most recent publicly available benchmark covering various VQA subtasks, not only SD questions. A summary of these results is presented in Table 5.1.
@@ 6 Configuration Analysis
Since our exploration algorithm relies on answer confidence, here we evaluate if the confidence values are well calibrated, i.e., the confidence correlates with the accuracy. For that, we adopt Reliability Diagrams [degroot1983comparison], which group predictions into bins based on confidence and measure the gap (calibration error) between confidence and accuracy for each bin. From these plots, we compute the Average Calibration Error (ACE) and Calibration Count (CC), quantifying the percentage of predictions above a defined confidence threshold, weighted by their accuracy calibration error (1-CE).
Given the poor performance of state-of-the-art VLMs on FALCON-Bench, we ease the task by providing directly one-minute clips taken from the middle of the ground-truth (GT) temporal interval (Sec. 10.1 of supp. material). For MCQs, confidence computation is straightforward, as only a single token is output (Sec. 3.2). However, for OQs, confidence must be aggregated across the entire sequence of tokens. Similar to calibration studies for LLMs [liu2023litcab], we investigate various aggregation metrics in Sec. 10.3, identifying the geometric average (Eq.1) as the most suitable approach for VLMs. Table 5 presents the calibration analysis for SOTA VLMs, with all methods showing strong calibration in MCQs. For OQs, Qwen2.5-VL significantly outperforms other approaches. The Section 10.3 of the supplementary material includes deeper calibration analysis, including the corresponding models’ calibration plots.
@@
We validate FALCONEye’s main features by first illustrating its training-free and model agnostic capabilities in Table 6, showing that current FALCONEye-Pro performance could be easily increased using a more powerful LLM or even VLM. Second, Table 7 shows the significance of using a calibrated confidence estimation of the VLM answers inside our exploration algorithm. To achieve this, we compare our confidence estimation mechanism against those used in related works: verbal confidence estimated together with the answer (by the VLM in our case) [zhi2025videoagent2] and verbal confidence estimated separately from the answer by an LLM given the question, answer and video context [wang2024videoagent]. We also perform a detailed ablation of the different FALCONEye’s exploration algorithm stages to show their contribution, including the zoom-in effect. The details of this study are in Sec. 10.2 of supplementary material.

### v3

@@ 4 FALCON-Bench
The videos of our benchmark were sourced from SoccerNet [20], MovieChat-1K [45] and Walking Tours Dataset [48], ensuring diversity and comprehensive coverage across different video types. Overall, the benchmark comprises 575 questions, covering 4 categories, over 90 videos, with an average video duration of 78.9 minutes and answers localized within a GT temporal window of 38.4 seconds. The dataset is split into a test set (506 questions) and a validation set (70 questions). Each question is accompanied by four different answer options, allowing for multiple-choice question (MCQs) evaluation. However, the primary focus remains on open-ended questions (OQs), as this approach is more aligned with the VAS task. The annotation process was conducted by humans assisted by GPT-4o for generating answer options. More details on the benchmark content and construction process are in Sec. 9 of the supplementary material. The benchmark has been built under the lmms-eval framework [63] to ease further evaluation and will be publicly available upon acceptance.
@@
mPLUG-Owl3 [58] 7B 16 5.52 41.1 28.5 36.0 35.2 0.00 6.63 8.82 4.23 10.0 7.68 0.68 0.29 0.66 0.54 0.00
LLaVA-v1.5 [28] 7B 32 11.4 23.5 31.9 44.0 33.1 0.00 12.7 3.92 4.80 2.00 3.57 0.36 0.31 0.24 0.30 0.00
LLaVA-v1.6 [23] 7B 32 12.9 28.4 30.5 38.0 32.3 0.00 14.3 4.90 4.51 2.00 3.80 0.31 0.30 0.20 0.27 0.00
LLaVA-OV [22] 7B 32 16.9 48.0 30.2 46.0 41.4 0.00 7.29 9.80 4.23 16.0 10.0 0.53 0.25 1.00 0.59 0.00
Qwen2.5-VL [4] 7B 768 68.9 50.0 24.4 48.0 40.8 0.00 70.1 11.7 8.19 10.0 9.96 0.72 0.53 0.62 0.62 0.00
@@
Video-LLaVA [26] 7B 8 3.37 28.4 26.8 34.0 29.7 0.00 4.50 6.86 5.08 8.00 6.64 0.56 0.37 0.68 0.53 0.00
VideoChat2-HD [25] 7B 8 6.74 29.4 24.2 26.0 26.5 0.00 8.06 8.82 2.82 8.00 6.54 0.71 0.20 0.52 0.47 0.00
LLaVA-Video [65] 7B 32 12.4 50.0 38.9 58.0 48.9 0.00 13.0 10.7 11.0 12.0 11.2 0.76 0.66 0.86 0.76 0.00
@@
MovieChat-OV [45] 7B 1024 62.3 43.1 38.7 46.0 42.6 0.00 69.3 9.80 4.80 10.0 8.20 0.53 0.25 0.66 0.48 0.00
Apollo [70] 7B 0.5fps 36.2 51.9 35.8 64.0 50.5 0.00 36.5 16.6 10.7 18.0 15.1 0.96 0.68 1.06 0.90 0.00
Video-XL-Pro [29] 3B 256 48.8 52.9 33.8 50.0 45.6 0.00 7B 17.6 7.62 14.0 13.0 1.08 0.45 0.94 0.63 0.00
LLaVA-OV+ReKV [15] 7B 0.5fps 92.1 65.6 40.1 56.0 53.9 0.00 96.6 34.3 13.8 24.0 24.0 1.78 0.80 1.40 1.32 0.00
@@
Socratic [61] 7B+N/A 0.5fps 186.6 53.9 48.0 54.0 51.9 21.3 186.7 25.4 29.0 20.0 24.8 1.49 1.71 1.16 1.45 19.6
LifelongMemory[53] 7B+N/A 0.5fps 207.3 45.0 41.5 38.0 41.8 0.00 208.2 6.86 22.3 8.00 12.3 0.53 1.37 0.56 0.82 0.00
VideoAgent[51]+OQ 7B+N/A 9.2 45.0 48.0 52.2 36.0 45.4 19.1 43.8 8.82 24.8 8.00 13.8 0.56 1.38 0.44 0.79 13.3
@@ 5 Experiments
In MCQs, we report accuracy as the primary evaluation metric. For OQs, following [31], also applied in others, as InfiniBench [3], we employ GPT-assisted evaluation. GPT reports accuracy (100 or 0) and a score value (0-5 range), assessing key aspects (Correctness of Information, Detailed Orientation, Contextual Understanding, Temporal Understanding and Consistency). We also use Ground Truth over Union (GToU) metric to evaluate the quality of the predicted temporal interval, as visual evidence (short clip) that contains the answer. Section 9.3 of the Supplementary discusses why we find GToU more adequate than mIoU (a clip that tries to match the entire GT clip).
@@
Socratic [61] 1 13.5K 2.1 12.6 768 14.6 186.7 24.8 1.45 19.6
Socratic-Short [61] 1 2.4K 1.1 12.6 64 9.8 125.0 15.5 1.03 18.3
LifelongMemory[53] 10.2 0.3K 1.0 12.6 64 9.8 208.2 12.3 0.82 0.00
VideoAgent[51]+OQ 5.6 1.3K 3.2 3.70 64 2.2 43.8 13.8 0.79 13.3
@@
A widely used baseline in recent long VQA benchmarks [6] is the Socratic method [61]. In this baseline, a VLM generates captions $\mathcal{C}=\{c_{1},\ldots,c_{n}\}$ from short video clips $\{v_{1},\ldots,v_{n}\}\in\mathcal{V}$ and a LLM generates $\mathcal{A}$ from $Q$ and $\mathcal{C}$. In addition to this baseline, we introduce three alternative exploration baselines (Fig. 7 of the supp. material) that rely solely on a VLM, guided by its answer confidence computed as described in Sec 3.
@@
As discussed in the related work, two relevant and recent video agents are VideoAgent [51] and VideoTree [54]. We use VideoAgent as our reference meta‑architecture to compare with since it couples CLIP and a 7B‑parameter VLM, both of which fit comfortably on a single RTX3090. In contrast, VideoTree depends on EVA‑CLIP 8B—too large to co‑load with a VLM‑7B on that GPU—and its public codebase lacks the VLM captioning module, preventing a fair, fully reproducible baseline.
@@
We evaluate our meta-architecture using Qwen2.5-VL [4] as the VLM and GPT4o-mini [35] as the LLM. As detailed in Sec. 6, Qwen2.5-VL outperforms other small-size VLMs on FALCON-Bench OQs in short clips near the answer, and is well calibrated. Its ViT supports images of any resolution, dynamically adjusting visual tokens, enabling an accurate zoom-in effect as clip lengths decrease. GPT4o-mini provides a balanced trade-off between cost and reasoning power (Sec. 6). Prompts used for LLM queries are detailed in Sec. 10.2. Inspired by the Socratic approach [6], first-level clips are 1min length, but captions are now shorter formed by up to 64 tokens (instead of 768). Within these clips, second-order (20s) and third-order (5s) clips are explored. The zoom-in effect, validated in supplementary Sec. 11, adjusts from 30 frames ($672\times 1204$) in first-level clips to 10 frames ($824\times 1462$) in third-order clips. In OQs, the LLM determines the final answer based on completion, confidence ($0.8$ threshold), captions, and temporal localization (as explained in Sec. 3). In MCQs, this evaluation is replaced by a simple confidence-based check with $0.9$ threshold since all possible answers are predefined (A, B, C, or D) and inherently valid, making confidence the only selection criterion. Lastly, we evaluate two versions of FALCONEye: FALCONEye-Pro, the top-performing configuration, which evaluates up to 45 candidate clips, and FALCONEye-Flash, a cost-efficient variant that limits evaluation to 10 candidates.
@@ 5.1 FALCON-Bench results
Table 4 presents the performance of baselines, open-source VLMs, and meta-architectures on FALCON-Bench. As baselines, we include human performance both to establish an upper bound and to analyze human strategies in VAS, which guided our algorithm design (Sec. 9.4). Additionally, we evaluate our two selected LLMs (GPT4o-mini and Qwen2.5-VL) in blind mode (without visual information) to assess their reliance on text-based priors. For open-source VLMs, we test small-scale versions of the main state-of-the-art models, including Apollo and MovieChat, which are specifically designed for long videos. We also evaluate three video agents, Socratic [61], LifelongMemory [53] for VQA and VideoAgent [51]. To ensure a fair comparison with our FALCONEye, these meta-architectures are implemented using the same VLM (Qwen2.5-VL) and LLM (GPT4o-mini). The Socratic approach generates captions (768 tokens) from 1-minute clips, while VideoAgent is extended to handle open-ended questions and predict answer localization, resulting in VideoAgent [51]++OQ. Most VLMs that are limited to a small number of sampled frames fail to outperform the LLM-blind baselines in MCQs, suggesting that their advantage over random guessing stems from discarding certain options, rather than actual answer retrieval from visual content. A similar issue arises with Qwen2.5-VL when operating in its default mode (2fps, up to 768 frames), as it downscales frames to $140\times 240$, losing crucial details. Only LLaVA-Video, Apollo, and ReKV show clear improvements over the blind baselines. This trend is not observed in OQs, where models must generate answers independently, validating their reliability. Among meta-architectures, only FALCONEye demonstrates robustness for long-form VAS, achieving 70.0% accuracy in MCQs and 44.7% in OQs with its top-performance configuration, and 64.7% and 41.1%, respectively, in the cost-efficient variant. The latter configuration achieves performance close to the Socratic baseline while maintaining a processing time per question not so far from human response times. We analyze their costs and performances, including exploration baselines, in Table 3. We also include Socratic-Short, which utilizes the short captions (64 tokens) generated by FALCONEye. While baseline exploration algorithms outperform existing meta-architectures, they require significantly more time. Notably, our LLM-guided exploration reduces this overhead while improving performance by efficiently narrowing the search.
@@ 5.2 FALCONEye generalization evaluation
To explore FALCONEye generalization beyond our benchmark, we evaluate it out-of-the-box on a standard VQA setting: shorter videos and a broader mix of question types beyond single-detail (SD). With this goal, we select MLVU [69], the most recent publicly available benchmark covering various VQA subtasks, not only SD questions. A summary of these results is presented in Table 5.1.
@@ 6 Configuration Analysis
Since our exploration algorithm relies on answer confidence, here we evaluate if the confidence values are well calibrated, i.e., the confidence correlates with the accuracy. For that, we adopt Reliability Diagrams [13], which group predictions into bins based on confidence and measure the gap (calibration error) between confidence and accuracy for each bin. From these plots, we compute the Average Calibration Error (ACE) and Calibration Count (CC), quantifying the percentage of predictions above a defined confidence threshold, weighted by their accuracy calibration error (1-CE).
Given the poor performance of state-of-the-art VLMs on FALCON-Bench, we ease the task by providing directly one-minute clips taken from the middle of the ground-truth (GT) temporal interval (Sec. 11.1 of supp. material). For MCQs, confidence computation is straightforward, as only a single token is output (Sec. 3.2). However, for OQs, confidence must be aggregated across the entire sequence of tokens. Similar to calibration studies for LLMs [30], we investigate various aggregation metrics in Sec. 11.3, identifying the geometric average (Eq.1) as the most suitable approach for VLMs. Table 5 presents the calibration analysis for SOTA VLMs, with all methods showing strong calibration in MCQs. For OQs, Qwen2.5-VL significantly outperforms other approaches. The Section 11.3 of the supplementary material includes deeper calibration analysis, including the corresponding models’ calibration plots.
@@
We validate FALCONEye’s main features by first illustrating its training-free and model agnostic capabilities in Table 6, showing that current FALCONEye-Pro performance could be easily increased using a more powerful LLM or even VLM. Second, Table 7 shows the significance of using a calibrated confidence estimation of the VLM answers inside our exploration algorithm. To achieve this, we compare our confidence estimation mechanism against those used in related works: verbal confidence estimated together with the answer (by the VLM in our case) [68] and verbal confidence estimated separately from the answer by an LLM given the question, answer and video context [51]. We also perform a detailed ablation of the different FALCONEye’s exploration algorithm stages to show their contribution, including the zoom-in effect. The details of this study are in Sec. 11.2 of supplementary material.
@@ 7 Conclusions
8 Acknowledgments
This work was supported by a DGA scholarship and by DGA project T45_23R, and grants AIA2025-163563-C31, PID2024-159284NB-I00, PID2021-125514NB-I00 and PID2024-158322OB-I00 funded by MCIN/AEI/10.13039/501100011033 and ERDF.